A university bookstore recently sold a wirebound graph-paper notebook for $1.10, and a college-ruled notebook for $3.16. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $102.38. How many of each type were sold?
Answer by rfer(16322)

c+g=50 g=50-c
1.10g+3.16c=102.38
1.10(50-c)+3.16c=102.38
55-1.10c+3.16c=102.38
55+2.06c=102.38
2.06c=102.38-55
2.06c=47.38
c=23 college-ruled
g=50-23
g=27 graph paper
-----------
27*1.10=$29.70
23*3.16=$72.68
total=$102.38